# Fibred sets within a predicative and constructive effective topos

### v1

@@ 1 Introduction
Furthermore, as already observed in [MM21], our constructions of $\mathbf{pEff}$ can be embedded in ${\mathbf{Eff}}$, and actually in its subcategory of the exact on lex completion of recursive objects, called discrete objects in [HRR90, vO15], thanks to the fact that ${\mathbf{Eff}}$ can be viewed as the exact on lex completion of the category of partitioned assemblies (see [RR90, MT22, MT21]).
It is worth recalling that a constructive predicative study of ${\mathbf{Eff}}$ was developed in the context of algebraic set theory by B. Van den Berg and I. Moerdijk (see [BM11]) by taking Aczel’s $\mathbf{CZF}$ in [AR01] as the set theory to be realized in such toposes. Our constructive and predicative $\mathbf{pEff}$ can be faithfully mapped in theirs, and more precisely in their subcategory of discrete objects [HRR90, vO15], when formalized in $\mathbf{CZF}$ + $\mathbf{REA}_{\bigcup}+\mathbf{RDC}$.
@@ 10 Comparison with related works
Now, the category $\mathcal{Eff}_{\mathcal{E}[\mathbf{T}]}$ proposed in [BM11] as a predicative version of Hyland’s Effective Topos ${\mathbf{Eff}}$ turns out to correspond exactly to the ex/reg completion of $\mathcal{Asm}_{\mathcal{E}[\mathbf{T}]}$, This follows from the fact that in $\mathbf{T}$ singletons are sets and that every diagonal in the ex/reg completion is covered by the image of a diagonal of assemblies.
@@
In addition, recalling that $\mathcal{Disc}_{\mathcal{E}[\mathbf{T}]}$ is the ex/reg completion of the category of modest sets (see [Ros90])), namely the category of assemblies which satisfy
@@
The category $\mathbf{pEff}$ built on an extension $\mathbf{T}$ of $\mathbf{CZF}$ having the numerical existence property is a subcategory of the predicative realizability category $\mathcal{Eff}_{\mathcal{E}[\mathbf{T}]}$ defined as in [BM11] from the syntactic category $\mathcal{E}[\mathbf{T}]$ of definable classes of $\mathbf{T}$ endowed with the family $\mathcal{S}[\mathbf{T}]$ of $\mathbf{T}$-provable set-fibered arrows.
@@
Observe that, when we consider $\mathbf{T}$ to be $\mathbf{ZFC}$, although it does not satisfy the numerical existence property, we can still define a functor from $\mathbf{pEff}$ to a version ${\mathbf{Eff}}_{def}$ of Hyland’s Effective Topos built on the category of definable sets of $\mathbf{ZFC}$. Indeed, since in $\mathbf{ZFC}$ all definable subclasses of $\mathbb{N}$ are in fact sets, one can define a faithful functor that is injective on objects from $\bm{\mathcal{C}}_{r}$ to an adequate category of recursive objects based on definable sets that cannot be proven to be full but still preserve the logical operators that are needed to show that the ex/lex completion of $\bm{\mathcal{C}}_{r}$ is a subcategory of the ex/lex completion of such a category of recursive objects. Since the latter embeds in ${\mathbf{Eff}}_{def}$ we obtain that $\mathbf{pEff}$ is a subcategory of ${\mathbf{Eff}}_{def}$. In the same framework, one can see that this ${\mathbf{Eff}}_{def}$ is a full subcategory of $\bf\mathcal{Eff}_{\mathcal{E}[\mathbf{ZFC}]}$.
${\mathbf{pEff}^{(\mathbf{ZFC})}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{notfull}$${{\mathbf{Eff}}_{def}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{Eff}_{\mathcal{E}[\mathbf{ZFC}]}}$

### v2

@@ 1 Introduction
Furthermore, as already observed in [MM21], our constructions of $\mathbf{pEff}$ can be mapped into ${\mathbf{Eff}}$, and actually in its subcategory of the exact on lex completion of recursive objects, called discrete objects in [HRR90, vO15], thanks to the fact that ${\mathbf{Eff}}$ can be viewed as the exact on lex completion of the category of partitioned assemblies (see [RR90, MT22, MT21]).
It is worth recalling that a constructive predicative study of ${\mathbf{Eff}}$ was developed in the context of algebraic set theory by B. van den Berg and I. Moerdijk (see [BM11]) by taking Aczel’s $\mathbf{CZF}$ in [AR01] as the set theory to be realized in such toposes. The full subcategory of $\mathbf{pEff}$ consisting of objects equipped small equivalence relations can be embedded into van den Berg and Moerdijk’s predicative version of ${\mathbf{Eff}}$, when formalized in $\mathbf{CZF}$ + $\mathbf{REA}_{\bigcup}+\mathbf{RDC}$. Instead, the whole constructive and predicative $\mathbf{pEff}$ can only be embedded into the category of discrete objects [HRR90, vO15] in the same metatheory.
@@ 10 Comparison with related works
In addition, recalling that $\mathcal{Disc}_{\mathcal{E}[\mathbf{T}]}$ is the ex/reg completion of the category of modest sets (see [Ros90]), namely the category of assemblies which satisfy
@@
Let $\mathbf{pEff}$ be the predicative effective topos built on an extension $\mathbf{T}$ of $\mathbf{CZF}$ having the numerical existence property; and let $\mathbf{pEff}_{sm.\Delta}$ be the full subcategory of $\mathbf{pEff}$ given by objects $(\mathsf{A},[\mathsf{R}])$ where $[\mathsf{R}]\in{\mathbf{Prop}^{r}_{\mathbf{s}}}(\mathsf{A}\times\mathsf{A})$. Then, $\mathbf{pEff}_{sm.\Delta}$ is a full subcategory of the predicative realizability category $\mathcal{Eff}_{\mathcal{E}[\mathbf{T}]}$ defined as in [BM11] from the syntactic category $\mathcal{E}[\mathbf{T}]$ of definable classes of $\mathbf{T}$ endowed with the family $\mathcal{S}[\mathbf{T}]$ of arrows whose fibers are sets in $\mathbf{T}$.
@@
Proof.
The category $\mathcal{Eff}_{\mathcal{E}[\mathbf{T}]}$ proposed in [BM11] as a predicative version of Hyland’s Effective Topos ${\mathbf{Eff}}$ is the full subcategory of the ex/reg completion of $\mathcal{Asm}_{\mathcal{E}[\mathbf{T}]}$, whose objects are separated in the sense of pg. 141 of [BM08]. The statement follows observing that objects $(\mathsf{A},[\mathsf{R}])$ of $\mathbf{pEff}_{sm.\Delta}$, seen as objects in $(\mathcal{Asm}_{\mathcal{E[\mathbf{T}]}})_{ex/reg}$, are separated in the above sense since $[\mathsf{R}]\in{\mathbf{Prop}^{r}_{\mathbf{s}}}(\mathsf{A}\times\mathsf{A})$. ∎
Observe that, when we consider $\mathbf{T}$ to be $\mathbf{ZFC}$, although it does not satisfy the numerical existence property, we can still define a functor from $\mathbf{pEff}$ to a version ${\mathbf{Eff}}_{def}$ of Hyland’s Effective Topos built as the ex/reg completion of the full subcategory of $\mathcal{Asm}_{\mathcal{E}[\mathbf{ZFC}]}$ whose objects are assemblies of which the underlying classes are definable sets. Indeed, in $\mathbf{ZFC}$ all definable subclasses of $\mathbb{N}$ are in fact sets. In the same framework, one can see that this ${\mathbf{Eff}}_{def}$ is a full subcategory of $\bf\mathcal{Eff}_{\mathcal{E}[\mathbf{ZFC}]}$.
${\mathbf{pEff}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{not\ full}$${{\mathbf{Eff}}_{def}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{Eff}_{\mathcal{E}[\mathbf{ZFC}]}}$